OpenStudy (anonymous):

find the x and y intercepts and graph the equation plotting the intercepts. Please show all of your work. -5x + 8y = -40

OpenStudy (bahrom7893):

x intercept: y = 0 -5x + 8*0 = -40 -5x = -40 x = 8 So: x intercept: (8;0)

OpenStudy (bahrom7893):

y intercept: x = 0 -5*0 + 8y = -40 8y = -40 y = -5

OpenStudy (anonymous):

you confirmed my answer, thanks!!

OpenStudy (bahrom7893):

np